OpenStudy (anonymous):

Put the following radical expression into simplified form. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\sqrt{\frac{ 48x^2y^3 }{ 7 }}\] Just can't seem to do these on my own correctly, ugh !

hartnn (hartnn):

i know you can do it on your own :) try it!

OpenStudy (anonymous):

I do, then it ends up wrong :/

hartnn (hartnn):

go step by step numerator : factor 48

OpenStudy (anonymous):

I got \[\frac{ 2xy^2\sqrt{21y} }{ 7 }\] That's incorrect...

hartnn (hartnn):

y^2 out ? shouldn't that be just y

hartnn (hartnn):

\(\sqrt {y^3} = \sqrt {y^2}\times \sqrt y \)

OpenStudy (anonymous):

16*3, 16=4 \[4\sqrt{3}\]

hartnn (hartnn):

4 sqrt 3 is correct so why did you put 2 instead of 4 ? :P

Miracrown (miracrown):

she mustn't been day dreaming - lol xD

hartnn (hartnn):

there are only 2 mistakes in your answer

OpenStudy (anonymous):

lol ugh, I can;t think straight right now lol

OpenStudy (anonymous):

i was doing fine earlier,smh

hartnn (hartnn):

first correction! make that 4 instead of 2 !

Miracrown (miracrown):

ah yes - uhm cookiibaii is it day at ur place?

OpenStudy (anonymous):

ehh, kind of lol it is morning, but it 3 am

hartnn (hartnn):

2nd correction \(\sqrt {y^3} = \sqrt {y^2}\times \sqrt y = y \sqrt y \)

Miracrown (miracrown):

there u go hartnn!

Miracrown (miracrown):

3am is the early start of the day, but overall to me 24 hours = 1 whole day. :P

OpenStudy (anonymous):

Got it! Thanks!